OpenStudy (anonymous):

use newton's method to obtain the third approximation, x2, of the positive fourth root of 12 by calculating the third approximation of the right 0 of f(x)= x^4 -12. start with x0=1

OpenStudy (anonymous):

so you want \[x^4-12=0\] first guess is 1 use \[x_2=x_1-\frac{f(x_1)}{f'(x_1)}\]

OpenStudy (anonymous):

across! where you been?

OpenStudy (anonymous):

\[f(1)=1^4-12=-11\] \[f'(x)=4x^3\] \[f'(1)=4\] \[x_2=1-\frac{-11}{4}=\frac{15}{4}=3.75\]

OpenStudy (anonymous):

next step is \[3.75-\frac{f(3.75)}{f'(3.75)}\] i would use a calculator